OpenStudy (anonymous):

I'm working through Session 1. I got the Sectant app to work and got all of the answers through it that I should have. My problem is that I can't get it to work out on paper. I'm going with the understanding that ΔY/ΔX should go like this ((.5x^3-x+ΔX)-x)/ΔX.

OpenStudy (anonymous):

Not quite. \[f \prime \left( x \right) = \lim_{\Delta x \rightarrow 0} \left[ \left( 0.5x^3 - x + \Delta x\right) - \left( 0.5x^3 - x) \right)\right]/\Delta x\] Hope this helps.

OpenStudy (anonymous):

The problem tells me that x = -0.75 and Δx= -0.5. When I run it [(0.5(-0.75)^3 −(-0.75)+-0.5)−(0.5(-0.75)^3 −(-0.75))]/-0.5 I get -2

OpenStudy (anonymous):

Okay, let's look at this a different way, so as not to get bogged down in complex calculations. The red point is (-0.75, 0.539) and the yellow point is (-1.25, 0.273). Taking \[\Delta y / \Delta x\] and rounding to the nearest hundredth gives the correct answer.

OpenStudy (anonymous):

Oh, and my original formula was wrong, sorry. The formula for the derivative is: \[\lim_{\Delta x \rightarrow 0} [f(x + \Delta x) - f(x)]/ \Delta x\]This works out as \[\lim_{\Delta x \rightarrow 0} [((x + \Delta x)^3 - (x + \Delta x)) - (x ^3 - x)] / \Delta x \]

OpenStudy (anonymous):

But remember, that formula is the derivative, which is not the same as \[\Delta y / \Delta x\]

OpenStudy (anonymous):

To summarize my somewhat incoherent solution, delta y over delta x is the change in y over the change in x. I recommend using a calculator instead of just hovering your mouse, you need more accuracy. You do not need the formula for derivatives because the derivative is not the same as delta y over delta x, although delta y over delta x may approximate the derivative. Also, to clarify, your derivative formula is a bit off. The formula for the derivative is \[\lim_{\Delta x \rightarrow 0} [f(x + \Delta x) - f(x)] / \Delta x\]which in this problem works out to be \[\lim_{\Delta x \rightarrow 0} [((x + \Delta x)^3 - (x + \Delta x)) - (x ^3 - x)]/ \Delta x\] and simplified makes \[\lim_{\Delta x \rightarrow 0} \Delta x^2 + 3x^2 + 3x \Delta x + 1 = 3x^2\] I hope this helps. Let me know if you have any questions about how I worked everything out.

OpenStudy (anonymous):

I've tried the formula you have given, using -.75 for x and -.5 for Δx, and still do not get the answer.

OpenStudy (anonymous):

Are you still using the formula for the derivative? Remember, the derivative is not the same as delta y over delta x. I just fixed your formula for future reference. What you need is delta y over delta x. Do you know how to find that?

OpenStudy (anonymous):

I guess I don't get any of what I'm trying to do. I can't get delta y over delta x and I'm not sure what I should be getting for an answer to the derivative.

OpenStudy (anonymous):

Okay, let's start from the beginning. delta y over delta x is equal to the change in y over the change in x. y = f(x), so y at the red point is equal to x^3 - x, which is (-0.75)^3 - (-0.75). y at the yellow point is equal to x^3 - x, which is (-1.25)^3 - (-1.25). Subtract the y at the red point from the y at the yellow point. Subtract the x at the red point from the x at the yellow point. Then take the y-difference over the x-difference. If we let the red point be (x1, y1) and the yellow point be (x2, y2), we have \[(y2 - y1) / (x2 - x1)\] I hope this helps, and I'm proud of you for trying to work it out on paper instead of just relying on the mathlet.

OpenStudy (anonymous):

Unless I'm completely crazy the answer should be 7.0625.

OpenStudy (anonymous):

This answer doesn't seem to match up with anything on the app or the solution sheet.

OpenStudy (anonymous):

Ok after taking a break and rechecking I finally got 0.53125. I need to get some sleep.